# A Hypernetwork-Based Approach to KAN Representation of Audio Signals

### v1

@@ Appendix A Extended ablation study
[24, 12, 6] 10,500 0.0007 37.5744 1.4904 12.6576 1.7192 0.8912 0.3229
[24, 24, 12, 6] 18,564 0.0006 38.4196 1.3828 13.5231 1.9078 0.9393 0.3847
[48, 12, 6] 19,908 0.0008 38.0106 1.3325 13.0259 2.1609 0.9384 0.3734
[48, 12, 12, 6] 21,924 0.0006 38.3097 1.3430 13.3983 2.1817 0.9517 0.3886
[48, 16, 8] 23,408 0.0007 38.2895 1.3119 13.3530 2.2985 0.9562 0.3756
[48, 24, 12] 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
[96, 48, 24] 102,480 0.0005 40.9599 1.1792 16.1224 3.5224 0.9898 0.1594
In comparison to the 1.5 s case, where a spline order of size 3 is clearly the optimal hyperparameter setting, its advantage is less evident in Table 17. KAN with a spline order equal to 5 achieves higher PNSR and SI-SNR scores, without falling too much behind in PESQ and STOI columns.
Table 17: KAN effectiveness in terms of diverse spline order sizes and 3 s long input
Spline Order Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
1 28,860 0.0006 38.5038 1.4711 13.6106 2.2903 0.9540 0.3299
2 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
3 33,300 0.0008 38.6639 1.2540 13.6820 2.5639 0.9704 0.3578
5 37,740 0.0006 39.0929 1.2663 14.1943 2.4499 0.9661 0.3376
7 42,180 0.0006 38.8824 1.3175 13.9657 2.1987 0.9573 0.3378
In table 18, table 19 and table 20, which correspond respectively to the ablations of grid size, learning rate, and enable scale spline flag on 3 s speeches, we observe similar conclusions and tendencies as in the 1.5s scenario. Described in Table 9 (Sect. 4.2), and Tables 10, 11(Appendix A), respectively.
Table 18: KAN effectiveness in terms of diverse grid size and 3 s long input
Grid size Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
1 11,100 0.0029 31.3359 2.0005 5.4555 1.3828 0.7652 0.3998
5 19,980 0.0007 38.1135 1.4668 13.2103 1.9651 0.9160 0.2399
10 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
12 35,520 0.0005 39.4903 1.2745 14.6419 2.6706 0.9739 0.3204
17 46,620 0.0006 39.7273 1.3224 14.8089 3.0683 0.9734 0.2646
Table 19: Effect of various learning rate values on 3 s recordings
Learning rate Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.01 31,080 0.0006 39.4961 1.3238 14.6358 2.5597 0.9692 0.3940
0.007 31,080 0.0006 39.2295 1.2804 14.3368 2.5669 0.9707 0.3685
0.006 31,080 0.0006 39.1037 1.2781 14.2218 2.5516 0.9700 0.3633
0.005 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
0.001 31,080 0.0007 37.7570 1.4674 12.8016 1.9302 0.9261 0.3020
Table 20: Enable scale spline flag ablation for 3 s audio source
Enable Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
False 28,860 0.0007 38.0986 1.4767 13.2310 2.1969 0.9360 0.4218
True 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
Finally, in Tables 21, 22, and 23 one can notice that the scale base and the scale spline hyperparameters do not significantly influence the performance. While manipulating scale base, all metrics except for CDPAM vary by less than 9% (the mentioned CDPAM under 16%). Moreover, apart from MSE that almost does not change, the changes in metric scores are not monotonic with respect to the changes in parameter values.
Table 21: Evaluation of different scale base factors on 3 s audio samples
Scale Base Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.25 31,080 0.0006 39.0743 1.2804 14.1621 2.4855 0.9654 0.3574
0.5 31,080 0.0006 39.1813 1.2744 14.2891 2.5813 0.9714 0.3485
1 31,080 0.0006 38.9647 1.2643 14.0714 2.5415 0.9693 0.3573
2 31,080 0.0006 38.8621 1.2980 13.9575 2.5120 0.9654 0.3932
4 31,080 0.0005 39.7455 1.3119 14.9256 2.3783 0.9612 0.4040
Table 22: Evaluation of different scale spline factors on 3 s audio samples
Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.25 31,080 0.0006 39.3700 1.2761 14.4961 2.6388 0.9713 0.3758
0.5 31,080 0.0006 38.9780 1.2698 14.0723 2.5493 0.9689 0.3678
1 31,080 0.0006 39.0889 1.2654 14.2169 2.4618 0.9643 0.3667
2 31,080 0.0006 38.9006 1.2735 13.9963 2.5288 0.9696 0.3663
4 31,080 0.0006 38.9421 1.2740 14.0168 2.4376 0.9666 0.3448
Table 23: Evaluation of different scale noise factors on 3 s audio samples
Scale Noise Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
0.4 31,080 0.0006 39.0755 1.2712 14.1628 2.5252 0.9659 0.3711
0.2 31,080 0.0006 39.1540 1.2631 14.2794 2.5684 0.9687 0.3604
0.1 31,080 0.0006 39.0749 1.2624 14.1858 2.5053 0.9698 0.3488
0.05 31,080 0.0006 39.0345 1.2699 14.1197 2.5075 0.9680 0.3601
0.025 31,080 0.0006 39.1100 1.2706 14.2064 2.5701 0.9679 0.3613

### v2

@@ A.2. INR
NERF (33,261) 3.6e-3 $\pm$ 4.5e-3 32.70 $\pm$ 5.12 1.21 $\pm$ 0.42 13.16 $\pm$ 5.05
SIREN (31,191) 3.7e-3 $\pm$ 2.5e-2 37.13 $\pm$ 5.18 1.10 $\pm$ 0.38 17.05 $\pm$ 8.37
FINER (31,191) 1.0e-3 $\pm$ 2.5e-3 40.89 $\pm$ 5.60 1.23 $\pm$ 0.53 21.36 $\pm$ 4.86
WIRE (31,191) 2.0e-3 $\pm$ 6.2e-3 37.13 $\pm$ 5.87 1.03 $\pm$ 0.38 17.50 $\pm$ 5.59
RFF (39,201) 4.9e-3 $\pm$ 2.1e-2 32.94 $\pm$ 4.86 1.25 $\pm$ 0.42 13.63 $\pm$ 4.89
KAN (33,768) 1.6e-3 $\pm$ 2.8e-3 36.98 $\pm$ 5.78 1.14 $\pm$ 0.45 17.46 $\pm$ 5.87
Table 19. Reconstruction metrics for UrbanSound8K. We report mean and standard deviation across 100 samples, 10 samples per split.
INR (# params) MSE PSNR LSD SI-SNR
@@
NERF (33,261) 3e-3 $\pm$ 4.9e-3 34.23 $\pm$ 5.35 1.25 $\pm$ 0.52 14.81 $\pm$ 6.33
SIREN (31,191) 1.8e-3 $\pm$ 4.8e-3 37.77 $\pm$ 5.79 1.14 $\pm$ 0.46 17.15 $\pm$ 12.58
FINER (31,191) 7.1e-4 $\pm$ 1.3e-3 40.46 $\pm$ 5.07 1.27 $\pm$ 0.52 20.98 $\pm$ 4.23
WIRE (31,191) 1.6e-3 $\pm$ 2.7e-3 38.28 $\pm$ 6.74 1.11 $\pm$ 0.45 18.65 $\pm$ 7.49
RFF (39,201) 4.1e-3 $\pm$ 7.2e-3 33.47 $\pm$ 6.05 1.23 $\pm$ 0.51 14.09 $\pm$ 7.41
KAN (33,768) 1.6e-3 $\pm$ 2.1e-3 36.92 $\pm$ 6.02 1.15 $\pm$ 0.49 17.37 $\pm$ 6.15
GenAI Usage Disclosure
The authors declare that generative AI tools were not used in the preparation of the manuscript or the development of the corresponding source code, apart from software tools for the evaluation of spelling and grammatical correctness.